# Dark Sector Tunneling Field Potentials for a Dark Big Bang

### v1

@@ Abstract
All of the significant evidence for dark matter observed thus far has been through its gravitational interactions. After 40 years of direct detection experiments, the parameter space for Weakly Interacting Massive Particles (WIMPs) as dark matter candidates is rapidly approaching the neutrino floor. In this light, we consider a dark sector that is strongly decoupled from the visible sector, interacting exclusively through gravity. In this model, proposed by Freese and Winkler [1], dark matter can be produced through a first-order phase transition in the dark sector dubbed “The Dark Big Bang”. In this study, we fully determine the allowed region of parameter space for the tunneling potential that leads to the realization of a Dark Big Bang and is consistent with all experimental bounds available.
Dark Sector; Cosmology; Early Universe; Phase Transitions
@@ I Introduction
Dark matter has been researched since the 1930s and the fact that it exists in the universe is well established. The early indicators of dark matter’s existence came from extragalactic data. To explain the velocities of galaxies within clusters [2] and the internal stability of galaxies themselves requires the presence of extra non-luminous mass [3, 4, 5]. Advanced measurements of the anisotropies in the Cosmic Microwave Background (CMB) have provided the strongest evidence for the existence of dark matter. In 2018, the Planck collaboration found that 27% of the energy of the universe is in dark matter, while only 5% is in Standard Model (SM) particles [6]. The interested reader can refer to this comprehensive history of the discovery and ongoing search for dark matter [7]. Modern research efforts aim to uncover the properties of dark matter particles through particle colliders [8], direct detection experiments [9], and observations of the universe [10, 11, 12].
WIMPs have been the leading candidate for particle dark matter since these are the lightest particles arising from a supersymmetric extension to the standard model. Supersymmetry is a theoretical model proposing the existence of an array of yet-undetected particles mirroring the standard model. Supersymmetry helps explain the mass value of the Higgs boson and the vacuum energy of the universe. Since WIMPs are the lightest of these particles, they should be the most easily detected. Yet besides a contentious signal from the DAMA collaboration (which may be verified or ruled-out in upcoming direct detection experiments such as COSINE-100 [13]), no WIMPs have been conclusively detected. In fact, the parameter space above the neutrino floor is almost ruled out entirely (see [14]). In this light, two of the possible paths for dark matter research are to come up with direct detection experiments below the neutrino floor [15] or to abandoned the weakly-interacting model entirely. This work will focus on the latter path.
Strongly decoupled dark matter only interacts with the visible sector through gravity. This is known as the “nightmare scenario” for physicists looking for dark matter. Gravity is by far the weakest of the fundamental forces in nature, so if this is the only way dark matter speaks to the visible universe, it will be very hard to detect. The Dark Big Bang theory, proposed in 2023 by Freese and Winkler [1], makes the nightmare a little less scary. If dark matter originated in a Dark Big Bang (DBB), gravity waves produced in the phase transition could be detected, providing a source of evidence for this decoupled dark matter [1].
@@
The theory of the Dark Big Bang is constrained to be consistent with standard $\Lambda$CDM cosmology and cosmological observations. The purpose of this paper is to impose these constraints onto the initial conditions of the DBB, explicitly determining the parameter space of the dark sector tunneling field potential which governs many aspects of the phase transition. We will show the parameter space for the tunneling field potential has two distinct regions. Parameters from one of the regions have not yet been utilized for a benchmark DBB, so this region will be of specific interest. We will show that a DBB with parameters chosen from this new region can produce the correct relic abundance of dark matter in the universe.
@@ II.1 Background
The Dark Big Bang occurs through a first-order phase transition in the dark sector. [^1] Other models considering higher-order phase transitions can be studied in the future [1]. Since the dark sector is initially cold, the phase transition occurs through the quantum tunneling of the $\phi$ field through a potential barrier (as opposed to thermal fluctuations causing $\phi$ to jump over the potential barrier). Initially, all of the dark sector energy density resides in the tunneling field $\phi$. This phase transition can be modeled like a particle in a potential with two non-degenerate minima. An illustrative example is shown in Fig.1[1].
Figure 1: Model tunneling field potential used for a first-order phase transition in the dark sector [1]. At the beginning of the universe, $\phi$ sits in the metastable minimum of the potential well. The discontinuous phase transition is realized when $\phi$ quantum tunnels through the potential barrier and eventually comes to rest in the ultrastable minimum. The height of the potential barrier is denoted by $V_{\rm{b}}$. The separation between the two minima is $\Delta\phi$. The change in potential of $\phi$ through the phase transition is $\Delta V$.
While the dark sector is in the false vacuum state ($\phi$ sits in the metastable minimum of $V(\phi)$), quantum fluctuations cause bubbles of the true vacuum to randomly nucleate throughout the universe. If the bubbles nucleate with a large enough radius, expansion will be energetically favored over collapse and the bubbles will expand and collide [16]. In Fig. 1, this can be modelled by the $\phi$ field quantum tunneling through the potential barrier and materializing at a height $V(\phi)=\Delta V$ near the ultrastable minimum with zero kinetic energy. As $\phi$ rolls towards the ultrastable minimum, the kinetic energy gained is stored in the expanding bubble walls [16]. $\phi$ will then undergo oscillations around the ultrastable minimum, which are dampened by couplings to dark sector fields. As $\phi$ comes to a rest in the true vacuum state, bubble walls collide, releasing their energy into dark sector fields and gravity waves. A first-order phase transition is phenomenologically similar to the phase transition of super-heated water to gas [16]: bubbles of the new phase (gas) nucleate and expand in the old phase (water) until all of the water has been turned into gas.
The Lagrangian for the dark sector is [1]
@@ II.2 Phase Transition Parameters
The subscript notation means the variable is evaluated at the time of the DBB. We have also defined the energy density of the dark sector before the phase transition as $\rho_{\rm{DS}}\simeq\rho_{\phi}=\Delta V$ [1]. The DBB theory accounts for the creation of all dark matter in the decay of the dark false vacuum, so dark matter particles must be produced efficiently in this phase transition. This requires the presence of light degrees of freedom in the dark sector $g_{\rm{DS}}$ that the bubble walls can always efficiently decay into [1]. With these light degrees of freedom in the dark sector, the end of the phase transition is considered a period of dark sector reheating, and the dark plasma produced quickly reaches a thermal equilibrium described by $T_{\rm{DS}}$ [1]. Gravity waves make up at most 5% of the dark sector energy density released in the phase transition [1], so the energy density in the dark sector after the phase transition can be approximated by [1]
$\rho_{\phi}=\Delta V\simeq\frac{\pi^{2}}{30}g_{\rm{DS}}(T_{\rm{DS}})T_{\rm{DS}}^{4}$ (8)
@@
where $g_{\rm{DS,*}}\equiv g_{\rm{DS}}(T_{\rm{DS,*}})$. Then $\alpha$ is the ratio of the dark sector reheating temperature to the visible sector temperature when the DBB occurs, up to their respective degrees of freedom
@@ II.3 Bounds From the Early Universe
The DBB is bounded to be consistent with cosmological observations and standard $\Lambda$CDM cosmology. The bounds are discussed at length in Ref. [1] and we will provide a brief overview here. The lower limit on $\alpha$ is determined by the requirement that the energy released in the phase transition can account for all dark matter in the universe today [1]. In mapping the power spectrum of CMB anisotropies, the Planck collaboration determined today’s dark matter density to be $\rho_{\rm{DM},0}=1.26$ keVcm-3  [6]. Using the scale factor at the time of the DBB $a_{*}$ to translate this to be a constrain on $\alpha$, the lower bound on the strength of the DBB becomes [1]
@@
The upper limit of $\alpha$ comes from existing bounds on the amount of extra energy density present in the early universe (in the DBB scenario, that extra energy density is $\rho_{\rm{DS}}$[1]. Extra energy density causes a speed up of the expansion of the universe realized as an increase in the Hubble parameter $H$. The speed up could cause weak interactions to freeze out earlier in the universe leading to an increase in the relativistic species populating the early universe $N_{\rm{eff}}$. $N_{\rm{eff}}$ counts the number of these relativistic species (not including photons), and is called the effective number of neutrinos (although particles other than neutrinos can contribute to $N_{\rm{eff}}$[17]. An increase in this number leads to dampening in the small scale (large $\ell$) power spectrum of CMB anisotropies. The Standard Model predicts $N_{\rm{eff}}=3.046$ [17] and the Planck collaboration has determined bounds on the change in this number $\Delta N_{\rm{eff}}=0.22\pm 0.15$ [6]. Taking the weaker bound $\Delta N_{\rm{eff}}<0.5$ leads to the following upper limit on $\alpha$ [1]
@@
The final constraint is on the latest a DBB can occur to be consistent with the $\Lambda$CDM picture of structure formation. Requiring that dark matter obtains the correct adiabatic perturbations to seed structure formation leads to the strongest constraint on the latest a DBB can occur [1]. $\rho_{\phi}$ is a smooth fluid at the time of the DBB, so the dark matter content must pick up perturbations from the visible sector radiation bath which dominates the universe [1]. The radiation bath exhibits adiabatic perturbations, which are realized in a common, local time shift of all background quantities in cosmological perturbation theory [17]. Adiabatic perturbations are defined by the equality $\delta\rho_{r}(\eta,\textbf{x})=\bar{\rho_{r}}^{\prime}\pi(\eta,\textbf{x})$ where $\eta$ is conformal time and x is a three dimensional space coordinate. This equation says that perturbed quantities at some spacetime point are the same as those in the background universe at a slightly shifted spacetime point. Adiabatic perturbations in visible sector radiation cause the DBB phase transition to occur at slightly different times throughout the universe. These time differences result in the perturbations becoming imprinted on the dark matter produced in the DBB (see [1]).
@@
The bounds on a DBB are shown in Fig. 2, which has been expanded from [1] to include higher temperatures. The main focus of our work is to map this parameter space onto the parameters determining the tunneling field potential. In the next section, we will show how the phase transition parameters can be recast as functions of the tunneling field parameters.
@@ III.1 Bubble Nucleation and Decay
where S is the Euclidean action of the bounce solution. The bounce solution comes from solving the Euclidean, or imaginary time ($t_{\rm{E}}=it$), equation of motion for $\phi$: $\Box_{\rm{E}}-V^{\prime}(\phi)=0$ with the boundary conditions $d\phi(0,\textbf{x})/dt_{E}=0$ and $\phi(\pm\infty,\textbf{x})=0$ [21]. In Euclidean time, $\phi$ starts in the metastable minimum at $\rm{t_{E}}=-\infty$, comes to rest near the ultrastable minimum at $\rm{t_{E}}=0$, and returns back to the metastable minimum at $\rm{t_{E}}=\infty$, hence the “bounce”. The action for the quartic potential considered here has a numerical solution beyond the analytic thin wall ($\Delta V\ll V_{\rm{b}}$) approximation [22]
$S\simeq\frac{\pi^{2}\mu^{6}}{24(\mu^{2}-2\lambda m^{2})^{3}}\left[13.832\frac{4\lambda m^{2}}{\mu^{2}}-10.819\left(\frac{4\lambda m^{2}}{\mu^{2}}\right)^{2}+2.0765\left(\frac{4\lambda m^{2}}{\mu^{2}}\right)^{3}\right]$ (15)
@@
Additional phase transition (PT) parameters such as the duration of the PT and the initial radius of the nucleated bubbles are shown in [1] but are not considered in this work.
@@ III.2 Mapping Equations
In order to map the parameter space for a DBB on to the input parameters for a tunneling field potential, the phase transition parameters need to be put in terms of $m$, $\mu$, and $\lambda$. First we will deal with $T_{*}$. One complication that is immediately apparent from Eq. 17 is the dependence on $g_{\rm{eff}}$, which also depends on $T_{*}$ [^2] This is only an issue when $g_{\rm{eff}}$ is a dynamic variable (below $T\sim 1.8\times 10^{5}$eV).. To accurately account for the degrees of freedom, we incorporate the degrees of freedom into our definition of temperature. Thus, once the the temperature scale of the DBB is calculated, the corresponding degrees of freedom can be easily determined and divided out. Combining Eqs. 171614, we find
$T_{*}g_{\rm{eff}}^{1/4}(m,\mu,\lambda)=\left(\frac{45M_{pl}^{2}}{2\pi^{2}}\right)^{1/4}\left(\frac{8\pi}{105}\right)^{1/8}\left(m^{4}\left(\frac{S(m,\mu,\lambda)}{2\pi}\right)^{2}e^{-S(m,\mu,\lambda)}\right)^{1/8}$ (18)
It is a simpler process putting $\alpha$ in terms of tunneling field parameters, since the degrees of freedom cancel out in the expansion. From Eqs. 7 and 18, $\alpha$ becomes
$\alpha(m,\mu,\lambda)\simeq\frac{60}{45M_{pl}^{2}}\left(\frac{105}{8\pi}\right)^{1/2}\left(m^{4}\left(\frac{S(m,\mu,\lambda)}{2\pi}\right)^{2}e^{-S(m,\mu,\lambda)}\right)^{-1/2}\Delta V(m,\mu,\lambda)$ (19)

### v3

@@ Abstract
All of the significant evidence for dark matter observed thus far has been through its gravitational interactions. After 40 years of direct detection experiments, the parameter space for Weakly Interacting Massive Particles (WIMPs) as dark matter candidates is rapidly approaching the neutrino floor. Moreover, there have been no signs of WIMPs in any particle production experiments so far. In this light, we consider a dark sector that is strongly decoupled from the visible sector, interacting exclusively through gravity. In this model, proposed by Freese and Winkler [1], dark matter can be produced through a first-order phase transition in the dark sector dubbed “The Dark Big Bang.” We fully determine the allowed region of parameter space for the tunneling potential that leads to the realization of a Dark Big Bang and is consistent with all experimental bounds available.
Dark Sector; Dark Matter; Cosmology; Early Universe; Phase Transitions
@@ I Introduction
Dark matter (DM) has been researched since the 1930s and the fact that it exists in the universe is well established. The early indicators of dark matter’s existence came from extragalactic data. To explain the velocities of galaxies within clusters [2] and the internal stability of galaxies themselves requires the presence of extra non-luminous mass [3, 4, 5]. Advanced measurements of the anisotropies in the Cosmic Microwave Background (CMB) have provided the strongest evidence for the existence of dark matter. In 2018, the Planck collaboration found that 27% of the energy of the universe is in dark matter, while only 5% is in Standard Model (SM) particles [6]. The interested reader can refer to this comprehensive history of the discovery and ongoing search for dark matter [7]. Modern research efforts aim to uncover the properties of dark matter particles through particle colliders [8], direct detection experiments [9], and observations of the universe [10, 11, 12].
WIMPs have been one of the leading candidates for particle dark matter since those hypothetical particles are naturally produced with the observed DM relic abundance. WIMPs exist naturally in any supersymmetric extension to the standard model of particle physics. Supersymmetry (SUSY) is a theoretical model proposing the existence of an array of yet-undetected particles mirroring the standard model. It helps explain the mass value of the Higgs boson and the vacuum energy of the universe. WIMPs, being the lightest of the non-SM particles required by SUSY, should be easily detectable in most models. Yet, besides a contentious signal from the DAMA collaboration [13, 14, 15], which may be verified or ruled-out in upcoming direct detection experiments such as COSINE-100 [16, 17], no WIMPs have been conclusively detected. In fact, the parameter space above the neutrino floor becomes more and more squeezed in view of the ever increasing sensitivity of direct detection experiments [e.g. 18]. Moreover, no SUSY signatures have been found by the Large Hadron Collider (LHC). In this light, exploring DM models outside of the DM paradigm becomes paramount.
Strongly decoupled dark matter only interacts with the visible sector through gravity. This is known as the “nightmare scenario” for physicists looking for dark matter. Gravity is by far the weakest of the fundamental forces in nature, so if this is the only way dark matter speaks to the visible universe, it will be very hard to detect. The Dark Big Bang theory has been proposed in 2023 by Freese and Winkler as an alternative to WIMPs. In this scenario, DM is generated in the early universe via a first order phase transition and is strongly decoupled. However, if dark matter originated in a Dark Big Bang (DBB), gravity waves (GWs) produced in the phase transition could be detected, providing a source of evidence for this decoupled dark matter [1]. For certain dark matter models, such as Dark WIMPs considered in this paper, evidence for dark matter produced in a DBB can be observed through correlated GW and CMB signals [1].
@@
By requiring that the Dark Big Bang model is consistent with standard $\Lambda$CDM cosmology and all available relevant cosmological observations, we fully determine the parameter space of the dark sector tunneling field potential. This is important, since it governs many aspects of the phase transition and thus the observable consequences (via GWs and CMB measurements) of the DBB. We will show the parameter space for the tunneling field potential has two distinct regions (see Fig. 5). Parameters from one of those two have not yet been utilized or explored in the original DBB work [1], so this newly uncovered region will be of specific interest. We will also explicitly show that a DBB with parameters chosen from this new region can produce the correct relic abundance of dark matter in the universe (see Fig. 9), and lead to GWs observable with upcoming Pulsar Timing Arrays (PTAs), such as IPTA and SKA (see Fig. 10).
The remainder of this paper is organized as follows: in Sec. II we review the main ingredients of the Dark Big Bang (DBB) model [1]; in Sec. III we describe the method we use in order to constrain the tunneling potential parameters that control the DBB phenomenology; in Sec. IV we find all regions in the tunneling potential parameter space that lead to a DBB and are consistent with all experimental bounds; in Sec. V we discuss the implications of the regions identified. We end with conclusions in Sec. VI and two Appendices.
@@ II.1 Background
The Dark Big Bang occurs through a first-order phase transition in the dark sector. [^1] Other models considering higher-order phase transitions can be studied in the future [1]. Since the dark sector is initially cold, the phase transition occurs through the quantum tunneling of the $\phi$ field through a potential barrier (as opposed to thermal fluctuations causing $\phi$ to jump over the potential barrier). Initially, all of the dark sector energy density resides in the tunneling field $\phi$. This phase transition can be modeled like a particle in a potential with two non-degenerate minima. An illustrative example is shown in Fig.1 (reproduced from [1]).
Figure 1: Model tunneling field potential used for a first-order phase transition in the dark sector (figure reproduced from [1]). At the beginning of the universe, $\phi$ sits in the metastable minimum of the potential well. The discontinuous phase transition is realized when $\phi$ quantum tunnels through the potential barrier and eventually comes to rest in the ultrastable minimum. The height of the potential barrier is denoted by $V_{\rm{b}}$. The separation between the two minima is $\Delta\phi$. The change in potential of $\phi$ through the phase transition is $\Delta V$.
While the dark sector is in the false vacuum state ($\phi$ sits in the metastable minimum of $V(\phi)$), quantum fluctuations cause bubbles of the true vacuum to randomly nucleate throughout the universe. If the bubbles nucleate with a large enough radius, expansion will be energetically favored over collapse and the bubbles will expand and collide [19]. In Fig. 1, this can be modelled by the $\phi$ field quantum tunneling through the potential barrier and materializing at a height $V(\phi)=\Delta V$ near the ultrastable minimum with zero kinetic energy. Once tunneled, the field $\phi$ rolls towards the ultrastable minimum, with its gained kinetic energy stored in the expanding bubble walls [19]. $\phi$ will then undergo oscillations around the ultrastable minimum, which are dampened by couplings to dark sector fields. As $\phi$ comes to a rest in the true vacuum state, bubble walls collide, releasing their energy into dark sector fields and gravity waves. A first-order phase transition is phenomenologically similar to the phase transition of super-heated water to gas [19]: bubbles of the new phase (gas) nucleate and expand in the old phase (water) until all of the water has been turned into gas.
The processes described above are controlled by the Lagrangian for the dark sector. Following [1], this is
@@ II.2 Phase Transition Parameters
The subscript notation means the variable is evaluated at the time of the DBB. We have also defined the energy density of the dark sector before the phase transition as $\rho_{\rm{DS}}\simeq\rho_{\phi}=\Delta V$ [1]. The DBB theory accounts for the creation of all dark matter in the decay of the dark false vacuum, so dark matter particles must be produced efficiently in this phase transition. This requires the presence of light (relativistic) degrees of freedom in the dark sector $g_{\rm{DS}}$ that the bubble walls can always efficiently decay into [1]. With these light degrees of freedom in the dark sector, the end of the phase transition is considered a period of dark sector reheating, and the dark plasma produced quickly reaches a thermal equilibrium described by $T_{\rm{DS}}$ [1]. Gravity waves make up at most 5% of the dark sector energy density released in the phase transition [1], so the energy density in the dark sector after the phase transition can be approximated by [1]
$\rho_{\phi}=\Delta V\simeq\frac{\pi^{2}}{30}g_{\rm{DS}}(T_{\rm{DS,*}})T_{\rm{DS,*}}^{4}$ (8)
@@
where $g_{\rm{DS,*}}\equiv g_{\rm{DS}}(T_{\rm{DS,*}})$. Then $\alpha$ can be written as the ratio of the dark sector reheating temperature to the visible sector temperature when the DBB occurs, up to their respective degrees of freedom
@@ II.3 Bounds From the Early Universe
The DBB is bounded to be consistent with cosmological observations and standard $\Lambda$CDM cosmology. The bounds are discussed at length in Ref. [1] and we will provide a brief overview here. The lower limit on $\alpha$ is determined by the requirement that the energy released in the phase transition can account for all dark matter in the universe today [1]. In mapping the power spectrum of CMB anisotropies, the Planck collaboration determined today’s dark matter density to be $\rho_{\rm{DM},0}=1.26$ keVcm-3  [6]. Requiring that all of the DM today was produced in a DBB leads to the following lower bound on the strength of the DBB (see [1])
@@
The upper limit of $\alpha$ comes from existing bounds on the amount of extra energy density present in the early universe (in the DBB scenario, that extra energy density is stored in $\rho_{\rm{DS}}$[1]. Extra energy density causes a speed up of the expansion of the universe realized as an increase in the Hubble parameter $H$. The speed up could cause weak interactions to freeze out earlier in the universe, leading to an increase in the relativistic species populating the early universe $N_{\rm{eff}}$. $N_{\rm{eff}}$ counts the number of these relativistic species (not including photons), and is called the effective number of neutrinos (although particles other than neutrinos can contribute to $N_{\rm{eff}}$[20]. An increase in this number leads to dampening in the small scale (large $\ell$) power spectrum of CMB anisotropies. The Standard Model predicts $N_{\rm{eff}}=3.046$ [20] and the Planck collaboration has determined bounds on the change in this number $\Delta N_{\rm{eff}}=0.22\pm 0.15$ [6]. This suggests $\Delta N_{\rm{eff}}<0.5$, which leads to the following upper limit on $\alpha$ [1]
@@
The final constraint is on the latest a DBB can occur to be consistent with the $\Lambda$CDM picture of structure formation. Requiring that dark matter obtains the correct adiabatic perturbations to seed structure formation leads to the strongest constraint on the latest a DBB can occur [1]. $\rho_{\phi}$ is a smooth fluid at the time of the DBB, so the dark matter content must pick up perturbations from the visible sector radiation bath which dominates the universe [1]. The radiation bath exhibits adiabatic perturbations, which are realized in a common, local time shift of all background quantities in cosmological perturbation theory [20]. Adiabatic perturbations are defined by the equality $\delta\rho_{r}(\eta,\textbf{x})=\bar{\rho_{r}}^{\prime}\pi(\eta,\textbf{x})$ where $\eta$ is conformal time and x is a three dimensional space coordinate. This equation says that perturbed quantities ($\delta\rho$) at some spacetime point are the same as those in the background universe ($\bar{\rho}$) at a slightly shifted spacetime point. Adiabatic perturbations in visible sector radiation cause the DBB phase transition to occur at slightly different times throughout the universe. These time differences result in the perturbations becoming imprinted on the dark matter produced in the DBB (see [1]).
@@
The bounds discussed above are summarized in Fig. 2, which has been expanded from [1] to include higher temperatures. The first aim of our work is to map this allowed parameter space (white gap region in Fig. 2) onto constraints on the parameters determining the tunneling field potential, which in turn control many aspects of a DBB model. In the next section, we will show how the phase transition parameters can be recast as functions of the tunneling field parameters.
@@ III.1 Bubble Nucleation and Decay
where S is the Euclidean action of the bounce solution. The bounce solution comes from solving the Euclidean, or imaginary time ($t_{\rm{E}}=it$), equation of motion for $\phi$: $\Box_{\rm{E}}-V^{\prime}(\phi)=0$ with the boundary conditions $d\phi(0,\textbf{x})/dt_{E}=0$ and $\phi(\pm\infty,\textbf{x})=0$ [24]. In Euclidean time, $\phi$ starts in the metastable minimum at $\rm{t_{E}}=-\infty$, comes to rest near the ultrastable minimum at $\rm{t_{E}}=0$, and returns back to the metastable minimum at $\rm{t_{E}}=\infty$, hence the “bounce”. The action for the quartic potential considered here has a numerical solution beyond the analytic thin wall ($\Delta V\ll V_{b}$) approximation [25]
$S\simeq\frac{\pi^{2}\mu^{6}}{24\lambda(\mu^{2}-2\lambda m^{2})^{3}}\left[13.832\frac{4\lambda m^{2}}{\mu^{2}}-10.819\left(\frac{4\lambda m^{2}}{\mu^{2}}\right)^{2}+2.0765\left(\frac{4\lambda m^{2}}{\mu^{2}}\right)^{3}\right]$ (15)
@@
where $M_{P}=2.44\times 10^{21}$ MeV is the reduced Planck mass. Additional phase transition (PT) parameters such as the duration of the PT and the initial radius of the nucleated bubbles are shown in [1] but are not considered in this work.
@@ III.2 Mapping Equations
In order to map the parameter space for a DBB on to the input parameters for a tunneling field potential, the phase transition parameters need to be put in terms of $m$, $\mu$, and $\lambda$. First we will deal with $T_{*}$. One complication that is immediately apparent from Eq. 17 is the dependence on $g_{\rm{eff}}$, which also depends on $T_{*}$. [^2] This is only an issue when $g_{\rm{eff}}$ is a dynamic variable (below $T\sim 1.8\times 10^{5}$eV). Combining Eqs. 171614, we find
$T_{*}g_{\rm{eff}}^{1/4}(m,\mu,\lambda)\simeq\left(\frac{45M_{P}^{2}}{2\pi^{2}}\right)^{1/4}\left(\frac{8\pi}{105}\right)^{1/8}\left(m^{4}\left(\frac{S(m,\mu,\lambda)}{2\pi}\right)^{2}e^{-S(m,\mu,\lambda)}\right)^{1/8}$ (18)
Note that for any given value of $T_{*}g_{\rm{eff}}^{1/4}$ one can uniquely solve for $T_{*}$, and consequently for $g_{\rm{eff}}(T_{*})$. It is a simpler process putting $\alpha$ in terms of tunneling field parameters, since the relativistic degrees of freedom in both the visible and dark sectors cancel out in the expansion. From Eqs. 7 and 18, $\alpha$ becomes
$\alpha(m,\mu,\lambda)\simeq\frac{60}{45M_{P}^{2}}\left(\frac{105}{8\pi}\right)^{1/2}\left(m^{4}\left(\frac{S(m,\mu,\lambda)}{2\pi}\right)^{2}e^{-S(m,\mu,\lambda)}\right)^{-1/2}\Delta V(m,\mu,\lambda)$ (19)